40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Mon Mar 25 02:04:05 2013, Koji, General, General, OMC building plan / procedure ~ WB18 
    Reply  Mon Mar 25 18:34:25 2013, Koji, General, General, OMC building plan / procedure ~ Mar 25 Mon 
       Reply  Wed Mar 27 20:54:45 2013, Koji, General, General, OMC building plan / procedure ~ Mar 26 Mon 
          Reply  Wed Mar 27 20:54:54 2013, Koji, General, General, OMC building plan / procedure ~ Mar 27 Wed 
             Reply  Wed Mar 27 20:55:10 2013, Koji, General, General, OMC building plan / procedure ~ Mar 28 Thu 
Message ID: 79     Entry time: Mon Mar 25 02:04:05 2013     Reply to this: 80
Author: Koji 
Type: General 
Category: General 
Subject: OMC building plan / procedure ~ WB18 

WB 18 March

  • Diode test
    • Dark current / Dark noise / Impedance
    • Quantum Efficiency test (but with glass)
    • Diode given to Bob for cleaning
  • Research possible issue of UV light on 1064 HR coating
    • ~ppm order loss increase after depositing 3J/cm^2 in 8 hours (i.e. same order to our illumination but in 10s for us)
    • Sent an e-mail to Ke-Xun Sun -> So far, no reply.
  • Glue test of PZT+prism+curved mirror with UV bond epoxy 
    • Done. Found some handling issues on the fixture.
  • Preparation of N2 line:
    • Done

Action items:

  • Bake test at 100°C for 1 hour at CIT 
    • Will be done on 25 Mon-26 Tue at Bob's lab
  • Curved mirror characterization
  • R&T measurement
ELOG V3.1.3-